OpenStudy (anonymous):

What are the period and phase shift for f(x) = 5 tan(2x − π)?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

is the period pi?

OpenStudy (anonymous):

@mathmate

Nnesha (nnesha):

y=Atan(Bx + C) + D a= amplitude for tan and cot period would be \[\rm P=\frac{ \pi}{\left| b \right|}\] phase shift c/b

OpenStudy (anonymous):

p=pi/2?

Nnesha (nnesha):

hmm no for sin and cos period would be 2pi/b for tan its pi/b

Nnesha (nnesha):

yes that's correct

OpenStudy (anonymous):

ok so the period would be pi/2 right

Nnesha (nnesha):

yes

OpenStudy (anonymous):

pi/2=phase shift?

Nnesha (nnesha):

hmm C is what ?

OpenStudy (anonymous):

What are the period and phase shift for f(x) = 5 tan(2x − π)? Period: π; phase shift: x = pi over two Period: π; phase shift: x = negative pi over two Period: pi over two; phase shift: x = negative pi over two Period: pi over two; phase shift: x = pi over two

Nnesha (nnesha):

okay so what is C in that equation ?

Nnesha (nnesha):

\[\rm y=Atan(Bx + C) + D\] f(x) = 5 tan(2x − π)?

OpenStudy (anonymous):

-pi/2 as the phase shift?

Nnesha (nnesha):

correct c is negative it's c= - pi

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

I have one last question.

OpenStudy (anonymous):

Use the graph below to answer the question that follows: cosine graph with points at 0, negative 1 and pi over 2, 3 and pi, negative 1 What are the amplitude, period, and midline of the function? Amplitude: 8; period: π; midline: y = 1 Amplitude: 8; period: 2π; midline: y = 5 Amplitude: 4; period: 2π; midline: y = 5 Amplitude: 2; period: π; midline: y = 1

Nnesha (nnesha):

\[\rm y=\color{green}{A}tan(\color{blue}{B}x + \color{Red}{C}) + \color{orange}{ D}\] \[\rm f(x) = \color{green}{5} \tan(\color{blue}{2}x \color{Red}{− π})+\color{orange}{0} \]

OpenStudy (anonymous):

ok. is the answer to the 2d question b?

Nnesha (nnesha):

where is the graph ? and tbh i'm not good at graphing these functions ..

OpenStudy (anonymous):

i cant copy it on here:O Just use desmos.com abd type it the points:)

Nnesha (nnesha):

take a scrreenshot like i mean how would i know what is -1 x or y ??

OpenStudy (anonymous):

@mathmate